OpenStudy (anonymous):

is cos^2 (x) equal to ( cosx )^2

OpenStudy (rational):

yes

OpenStudy (rational):

\(\large \cos^2(x) = (\cos(x) )^2\)

OpenStudy (rational):

however, \(\large \cos(x^2)\) is a different thingy